OpenStudy (1018):

please help me integrate this :

OpenStudy (1018):

\[56x^{13} e^{-x^{14}}\]

OpenStudy (freckles):

did you try a subsitution

OpenStudy (freckles):

let u=-x^(14)

OpenStudy (1018):

yeah it says i should use substitution rule but im still confused.

OpenStudy (freckles):

well if u=-x^(14) then du/dx=?

OpenStudy (1018):

i already have the answer, but i want to know how to get there

OpenStudy (freckles):

\[u=-x^{14} \\ \frac{du}{dx}=?\]

OpenStudy (freckles):

you know power rule

OpenStudy (freckles):

use it to find the derivative of -x^(14)

OpenStudy (1018):

is it -x^15 / 15 ?

OpenStudy (freckles):

derivative of -x^(14)?

OpenStudy (1018):

oh wait sorry thats wrong

OpenStudy (1018):

yeah sorry i saw it wait im gonna give you another answer

OpenStudy (1018):

-14x^13 hows that?

OpenStudy (freckles):

\[(c \cdot x^{n})'=c (x^{n})'=c \cdot n x^{n-1}\] this is power rule for differentiating yes that!

OpenStudy (freckles):

\[u=-x^{14} \\ \frac{du}{dx}=-14 x^{13} \\ du=-14x^{13} dx \\ -4 du=56 x^{13} dx\]

OpenStudy (freckles):

can you finish now?

OpenStudy (1018):

yes thank you!